OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. 2x + y + z = -7 x - 3y + 4z = -14 x - 2y - 3z = -11

OpenStudy (radar):

The method that I would choose is the "elimination" method. This may not have been the method you or others would choose. Add the last two equations getting: 2x -5y + z = -25. Now subtract this equation from the first original equation: 2x + y + z = -7 -) 2x -5y +z = -25 ------------------ 6y = 18 or y = 3 this has solved the problem to one of its variables. I chose the elimination method because it is the simplest for me.

OpenStudy (anonymous):

So is the correct answer B @radar

OpenStudy (anonymous):

Nvm